### v2

# Langevin equations and a geometric integration scheme for the overdamped limit of homogeneous rotational Brownian motion

Abstract
The translational motion of anisotropic or self-propelled colloidal particles is closely linked with the particle’s orientation and its rotational Brownian motion. In the overdamped limit, the stochastic evolution of the orientation vector follows a diffusion process on the unit sphere and is characterised by an orientation-dependent (“multiplicative”) noise. As a consequence, the corresponding Langevin equation attains different forms depending on whether Itō’s or Stratonovich’s stochastic calculus is used. We clarify that both forms are equivalent and derive them from a geometric construction of Brownian motion on the unit sphere, based on infinitesimal random rotations. Our approach suggests further a geometric integration scheme for rotational Brownian motion, which preserves the normalisation constraint of the orientation vector exactly. We show that a simple implementation of the scheme, based on Gaussian random rotations, converges weakly at order 1 of the integration time step, and we outline an advanced variant of the scheme that is weakly exact for an arbitrarily large time step. Due to a favourable prefactor of the discretisation error, already the Gaussian scheme allows for integration time steps that are one order of magnitude larger compared to a commonly used, ad hoc algorithm for rotational Brownian dynamics simulations. The discussion is restricted to time-homogeneous rotational Brownian motion (i.e., a single rotational diffusion constant), which is relevant for axisymmetric particles and also chemically anisotropic spheres such as self-propelled Janus particles.
I Introduction
The past two decades have seen considerable advances in the synthesis and characterisation of anisotropic colloidal particles, ranging from complex shapes, e.g., ellipsoids [1, 2], clusters [3], and propellers [4], to anisotropic chemical surface coatings. An important example for the latter are so-called Janus spheres [5, 6, 7, 8, 9, 10], which exhibit autonomous self-propulsion under suitable conditions and often serve as colloidal models for microswimmers. The use of such particles in experimental model systems has opened new avenues and has boosted research in the field of active matter (see, e.g., Refs. [11, 12, 13]). These successes have stimulated a large number of theoretical investigations, which have brought forth elaborate descriptions of the self-propelled motion of a single active colloid (see, e.g., Refs. [14, 15, 16, 17, 18, 19]). Depending on the question at hand, hydrodynamic effects due to solvent flows can be crucial [20, 21, 22, 23] or may be negligible. In the latter case, the active Brownian particle (ABP) model is widely used to describe the motion of active colloids [24, 17, 25]; it combines the Brownian motion of a colloidal Janus sphere with a constant propulsion velocity which randomly changes its direction. The propulsion direction is described by a unit vector $\boldsymbol{u}$ which itself performs a Brownian motion on the unit sphere. Thereby, the stochastic motion of $\boldsymbol{u}$ may evolve freely [5, 26, 27] or under the influence of gravity [28, 29, 30, 31, 32, 33] and it becomes particularly challenging in the case of shape anisotropy [34, *Doi:2005, 36, 37, 38, 39, 40, 41, *Leitmann:2017, 43]. The dynamics of active colloids becomes complex in the presence of rotation–translation coupling, e.g., due to short-time hydrodynamic friction [34, *Doi:2005, 36, 37] or due to confinement by a harmonic potential well [44] or a substrate potential [27, 45]; in such situations, the description of the rotational dynamics is crucial for predictions on the translational motion.
Models for the rotational motion of molecules have traditionally been studied in the context of dynamic light scattering and dielectric spectroscopy [46, 47]. These models are typically based on the Fokker–Planck equation for orientational diffusion [48, 49, 50], which allows one to calculate correlation functions and, more generally, the statistics of the trajectories [40, 41, *Leitmann:2017, 43]. Alternatively, some form of Langevin equation is employed, which provides a description on the level of stochastic trajectories and which forms the basis of stochastic simulations. The underdamped Langevin equation is used when inertia effects are relevant [46] and, mathematically, there are no conceptual questions with this equation apart from being demanding to work with it analytically. In the context of colloidal motion, however, one would like to model the stochastic dynamics as (completely) overdamped, which requires a suitable Wiener process for the particle orientation. The corresponding Langevin equation contains a noise term that has a “multiplicative” structure [46, 34, *Doi:2005, 36, 37, 29, 31, 25, 41, *Leitmann:2017, 43, 27, 45] [see below, Eqs. 7 and 9], which brings up questions about the underlying stochastic calculus. In particular, in its Itō form, the overdamped Langevin equation [Eq. 9] contains a drift-like term, which may appear counter-intuitive at first sight. Another example for this mathematically well-understood situation [51, 52] is given by translational diffusion near a wall characterised by a position-dependent diffusivity, which has lead to a re-arising apparent controversy over decades [53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. The issue of multiplicative noise in rotational diffusion is entirely circumvented by modelling the particle orientation as confined to a plane, which allows one to parametrise the two-dimensional rotations by a single angle. Whereas this choice is widely adopted in models of active matter, it often appears to be a choice of convenience for simplicity rather than being physically well justified. Notably, the linear transport behavior can sensitively depend on the dimensionality of rotational diffusion [45].
In this note, we aim to clarify some mathematical aspects of rotational Brownian motion, in particular, the different forms of the overdamped Langevin equation and how they depend on the adopted definition of the stochastic integral (Itō or Stratonovich). The equations are derived from a geometric construction of Brownian motion on the unit sphere, given as a sequence of infinitesimal random rotations. The construction suggests a geometric integration scheme for rotational Brownian motion, which (i) exactly preserves the normalisation constraint of the orientation vector and (ii) can be extended such that it is weakly exact for arbitrarily large time steps. In order to focus on the essence of the problem, we shall mainly discuss chemically anisotropic spheres (e.g., Janus spheres), thereby avoiding additional difficulties that arise from the incessant transformation of the diffusivity tensor due to the rotational motion. We then show that the results apply also to axisymmetric particles, however, their derivation requires some modifications which are mostly of technical nature. We refer to a Brownian motion with constant coefficients as homogeneous in time.
Finally, we note that the literature describes a number of numerical schemes for Brownian motion of rigid bodies, which are often justified only heuristically; this includes schemes using finite rotations [63, 64], adding a constraint force [65, 66, 67, 68], discretising the Stratonovich–Langevin equation [34, 69], or employing a symplectic splitting scheme of the underdamped Langevin dynamics [70]. Also the algorithm presented below (Section IV.1) has already been used in previous work of both authors [41, *Leitmann:2017, 27, 45], without further explanation.
II Heuristic derivation of the overdamped limit
II.1 Langevin equation of rotational Brownian motion with inertia
Let us start from classical mechanics and consider the rotational motion of a rigid body in three-dimensional space, which may be subject to a total torque $\boldsymbol{T}(t)$ at time $t$. Working in a space-fixed (or inertial) frame of reference, the body’s angular velocity $\boldsymbol{\omega}(t)$, obeys Euler’s equation, which is the rotational analogue of Newton’s law:
$\frac{\mathrm{d}}{\mathrm{d}t}[\boldsymbol{I}(t)\,\boldsymbol{\omega}(t)]=\boldsymbol{T}(t)\,;$ (1)
here, $\boldsymbol{I}(t)$ is the body’s moment of inertia, which depends on its instantaneous orientation in space. We consider a body-fixed axis, which is denoted by the unit vector $\boldsymbol{u}(t)$. Its time evolution given $\boldsymbol{\omega}(t)$ obeys the kinematic equation
$\dot{\boldsymbol{u}}(t)=\boldsymbol{\omega}(t)\times\boldsymbol{u}(t)\,.$ (2)
One readily verifies that this dynamics preserves the norm $|\boldsymbol{u}(t)|=1$, since $\mathrm{d}|\boldsymbol{u}|^{2}/\mathrm{d}t=2\boldsymbol{u}\cdot\dot{\boldsymbol{u}}=0$.
If the particle is immersed in a fluid medium, it experiences a torque $-\boldsymbol{\zeta}_{R}(t)\,\boldsymbol{\omega}(t)$ due to viscous (Stokes) friction. In general, the rotational friction coefficient $\boldsymbol{\zeta}_{R}(t)$ is a tensor which is constant in a body-fixed frame [71], but is time-dependent in the space-fixed frame used here. Following Langevin’s and Ornstein’s approach to Brownian motion, the incessant collisions of the particle with solvent molecules are further subsumed in a random torque $\boldsymbol{\xi}(t)$ acting on the particle. Taking the collisions to be uncorrelated in a first approximation suggests to model $\boldsymbol{\xi}(t)$ as a Gaussian white noise process, which is independent of $\boldsymbol{\omega}(s)$ for $t>s$, has mean zero, and its covariance matrix is determined by the fluctuation–dissipation relation:
$\langle\boldsymbol{\xi}(t)\rangle=0\quad\text{and}\quad\langle\boldsymbol{\xi}(t)\otimes\boldsymbol{\xi}(t^{\prime})\rangle=2k_{\text{B}}T\boldsymbol{\zeta}_{R}(t)\delta(t-t^{\prime})\,,$ (3)
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}[\boldsymbol{I}(t)\,\boldsymbol{\omega}(t)]=\boldsymbol{T}_{\text{ext}}(t)-\boldsymbol{\zeta}_{R}(t)\,\boldsymbol{\omega}(t)+\boldsymbol{\xi}(t)\,.$ (4)
This is a stochastic differential equation in $\boldsymbol{\omega}(t)$ and since the white noise $\boldsymbol{\xi}(t)$ enters “additively” (i.e., it is not multiplied by a $\boldsymbol{\omega}(t)$-dependent function), there is no ambiguity about the choice of the stochastic calculus: any (consistent) definition of the stochastic integral yields the same solution [52].
For the purpose of keeping the present discussion focused, we will limit ourselves to spherical particles with a chemically patterned surface so that their orientation remains identifiable; colloidal Janus particles are a prominent example. In this case, the friction tensor is proportional to the unit matrix and, thus, becomes time-independent: $\boldsymbol{\zeta}_{R}(t)=\zeta_{R}\mathds{1}$. Correspondingly, the moment of inertia reduces to a constant scalar as well: $\boldsymbol{I}(t)=I\mathds{1}$. If the sphere has the mass $m$ and the radius $a$ and is suspended in a solvent of dynamic viscosity $\eta$ and density $\varrho$, one finds $I=(2/5)ma^{2}$ and the solution of the Stokes flow yields $\zeta_{R}=8\pi\eta a^{3}$ [71].
In Section V, we extend the results to axisymmetric particles, for which $\boldsymbol{u}(t)$ is naturally chosen to align with the symmetry axis of the particle. This axis fully determines the particle’s orientation such that the friction tensor $\boldsymbol{\zeta}_{R}(t)$ and moment of inertia $\boldsymbol{I}(t)$ are functions of $\boldsymbol{u}(t)$. For particles of arbitrary shape, the orientation may be tracked by following the motion of a body-fixed trihedron [34, *Doi:2005, 36], an elaboration on this generalisation is left for future research.
II.2 Heuristic derivation of the overdamped limit
[…]
$\zeta_{R}\,\boldsymbol{\omega}(t)=\boldsymbol{T}_{\text{ext}}(t)+\boldsymbol{\xi}(t)\,.$ (5)
This simple reasoning is possible since Eq. 4 is of the Ornstein–Uhlenbeck type, in particular, since it is linear in $\boldsymbol{\omega}(t)$ and the noise $\boldsymbol{\xi}(t)$ enters additively. The same result would be obtained from the formal solution for $\boldsymbol{\omega}(t)$, given in integral form [52], and taking the limit $\tau_{1}\to 0$.
Equation 5 states that the angular velocity $\boldsymbol{\omega}(t)$ is statistically equivalent to a Gaussian white noise process with a deterministic bias $\boldsymbol{T}_{\text{ext}}(t)$. More precisely, we may take $\boldsymbol{\omega}(t)$ to be a Gaussian process which satisfies
$\langle\boldsymbol{\omega}(t)\rangle$ $=\zeta_{R}^{-1}\boldsymbol{T}_{\text{ext}}(t)\,,$
$\langle\delta\boldsymbol{\omega}(t^{\prime})\otimes\delta\boldsymbol{\omega}(t)\rangle$ $=2D_{\text{R}}\boldsymbol{1}\delta(t^{\prime}-t)$ (6)
[…]
It is now natural to interpret the kinematic equation (2) for the orientation $\boldsymbol{u}(t)$ as a stochastic differential equation with a random angular velocity $\boldsymbol{\omega}(t)$. We repeat the equation here:
[…]
which is one form of the overdamped Langevin equation for rotational Brownian motion of a body-fixed axis $\boldsymbol{u}(t)$. However, the white noise $\boldsymbol{\omega}(t)$ multiplies the variable $\boldsymbol{u}(t)$ and the re-interpretation of Eq. 2 is yet ill-defined unless we specify the stochastic calculus  [^1] We emphasise that we have left open this question when solving Eq. 4., i.e., unless we give a meaning to the time integral $\int_{0}^{t}\boldsymbol{\omega}(s)\times\boldsymbol{u}(s)\,\mathrm{d}s$. This is a modelling decision. Recalling that $\boldsymbol{u}(t)$ is a unit vector, Eq. 7 describes a diffusion process on a constraining manifold, and the natural choice in this situation is to interpret the time integral as a (Fisk–)Stratonovich integral (see chap. 35 in ref. 79).
In the physics literature, the Stratonovich interpretation of Eq. 7 is often (tacitly and correctly) assumed. For the numerical integration, however, a direct application of the widespread Euler–Maruyama algorithm, which is a simple and robust algorithm for Itō SDEs, would break the conservation of the norm $|\boldsymbol{u}(t)|$. This could be fixed a posteriori by rescaling $\boldsymbol{u}(t)$ after every integration step [66], but such an approach appears conceptually unsatisfying  [^2] In fact, this debatable ad hoc algorithm is currently implemented in the molecular dynamics simulation software LAMMPS (since release from 29 September 2021), see the command “fix Brownian/sphere” [see https://docs.lammps.org/fix_brownian.html]. . Instead, one should use a suitable discretisation scheme for the Stratonovich integral, e.g., the stochastic Milstein algorithm; yet, the correct implementation of such schemes is more demanding [81, 68]. A conceptually less satisfying alternative is to mix different stochastic calculi: one may cast the Stratonovich form (7) into an equivalent form that invokes the Itō integral and which is thus suitable for the Euler–Maruyama scheme [82, 46, 52]; this form includes a noise-induced term appearing as a drift [see below, Eq. 9]. We note that all of these integration schemes preserve the normalisation of $\boldsymbol{u}(t)$ only asymptotically for a vanishing integration time step, $\Delta t\to 0$.
II.3 Two-dimensional case.
For motion in two dimensional space, which is widely considered in the theoretical literature on active particles, the issue of the multiplicative noise can be circumvented since the angular velocity has only a single component. Writing $\boldsymbol{\omega}=(0,0,\omega)$, the component $u_{z}$ of $\boldsymbol{u}=(u_{x},u_{y},u_{z})$ remains unchanged under the dynamics so that we can set $u_{z}=0$ and parametrise the remaining two components $(u_{x},u_{y})=(\cos(\varphi),\sin(\varphi))$ in terms of the angle $\varphi\in[-\pi,\pi)$. The evolution of $\varphi(t)$ follows from Eq. 2, which reads for the first vector component: $\dot{u}_{x}(t)=-\omega\,u_{y}(t).$ Combining with the ordinary chain rule, $\dot{u}_{x}(t)=-\sin(\varphi(t))\,\dot{\varphi}(t),$ yields:
$\dot{\varphi}(t)=\omega(t)\,.$ (8)
Passing to the overdamped limit, we replace $\omega(t)$ with the white noise given in Eq. 5, which enters additively. In fact, Eq. 8 is the simplest form of a stochastic differential equation, which has an unambiguous solution with any stochastic calculus. We conclude that $\varphi(t)$ is a Brownian motion on the one-dimensional torus $[-\pi,\pi)$, i.e., $\boldsymbol{u}(t)$ performs a Brownian motion on the unit circle in the $xy$-plane. The same result would be obtained when starting from Eq. 7 and observing that the classical (Leibniz) chain rule still applies for the Stratonovich calculus.
III Geometric derivation of the overdamped Langevin equation
For analytical and numerical work the use of Itō’s calculus has certain advantages and, to this end, Eq. 7 needs to be cast into its Itō form, which reads:
[…]
with $\tau_{\text{R}}^{-1}:=(d-1)D_{\text{R}}$ for motion in $d$-dimensional space ($d=2,3$). The transformation is straightforward [52], but a bit tedious due to the vector cross product on the r.h.s. of Eq. 7 and will be given below, after an alternative derivation which is based on geometric considerations. We will also show that, due to the specific rules of Itō calculus, the additional term $-\tau_{\text{R}}^{-1}\boldsymbol{u}(t)$ does not generate a drift, but is necessary to preserve the norm $|\boldsymbol{u}(t)|$ and to yield the orientational autocorrelation function correctly. Furthermore, the solution $\boldsymbol{u}(t)$ is non-anticipating with respect to the noise $\boldsymbol{\omega}(t)$, i.e., the (Itō) increments of the noise, $\int_{t}^{t+\Delta t}\boldsymbol{\omega}(s)\mathrm{d}s$ are independent of $\boldsymbol{u}(s)$ for any $t\geqslant s$ and $\Delta t>0$. (This central statement has no simple analogue for the Stratonovich integral.) As a consequence, the Itō–Langevin equation (9) is amenable to the straightforward Euler–Maruyama scheme for numerical integration [81]. Below, we propose a geometric integration scheme that satisfies the constraint $|\boldsymbol{u}(t)|=1$ exactly for arbitrary $\Delta t$.
III.1 Geometric derivation of the Itō form
In the following, we adopt a geometric perspective to derive Eq. 9. For Brownian motion of the vector $\boldsymbol{u}(t)$ on the three-dimensional unit sphere, the evolution of $\boldsymbol{u}(t)$ emerges from the repeated action of infinitesimal random rotations, which are the analogue of the random displacements of planar Brownian motion. Thus, we can write:
$\boldsymbol{u}(t+\mathrm{d}t)=\text{e}^{\boldsymbol{\omega}(t)\mathrm{d}t\cdot\boldsymbol{J}}\boldsymbol{u}(t)\,,$ (10)
where $|\boldsymbol{\omega}(t)|\mathrm{d}t$ and $\boldsymbol{\omega}(t)/|\boldsymbol{\omega}(t)|$ are the angle and the axis of rotation at time $t$, respectively. The symbol $\boldsymbol{J}=(J_{1},\dots,J_{3})$ denotes the $3\times 3$ matrices $(J_{i})_{jk}=-\varepsilon_{ijk}$, given in terms of the Levi–Civita symbol. These matrices form a basis of the Lie algebra $\mathfrak{so}(3)$ such that $\boldsymbol{\omega}\cdot\boldsymbol{J}=\omega_{i}J_{i}$ is the skew-symmetric matrix representation of the axial vector $\boldsymbol{\omega}$, where summation of repeated indices is implied; written explicitly in components $\boldsymbol{\omega}=(\omega_{x},\omega_{y},\omega_{z})$:
$\boldsymbol{\omega}\cdot\boldsymbol{J}=\begin{pmatrix}0&-\omega_{z}&\omega_{y}\\ \omega_{z}&0&-\omega_{x}\\ -\omega_{y}&\omega_{x}&0\end{pmatrix}$ (11)
In particular, it holds
$\boldsymbol{\omega}\times\boldsymbol{v}=(\boldsymbol{\omega}\cdot\boldsymbol{J})\boldsymbol{v}$ (12)
for any vectors $\boldsymbol{\omega}$ and $\boldsymbol{v}$. The ansatz (10) resembles the McKean–Gangolli injection[83] of a stochastic process, here from the Lie algebra $\mathfrak{so}(3)$ into its Lie group $SO(3)$.
Expanding the exponential in Eq. 10 up to the order of $[\boldsymbol{\omega}(t)\mathrm{d}t]^{2}$ yields, according to the rules of Itō calculus,
$\boldsymbol{u}(t+\mathrm{d}t)$ $=\left[\mathds{1}+\boldsymbol{\omega}(t)\mathrm{d}t\cdot\boldsymbol{J}+\tfrac{1}{2}(\boldsymbol{\omega}(t)\mathrm{d}t\cdot\boldsymbol{J})^{2}\right]\boldsymbol{u}(t)$
$=\boldsymbol{u}(t)+\boldsymbol{\omega}(t)\mathrm{d}t\times\boldsymbol{u}(t)-2D_{\text{R}}\boldsymbol{u}(t)\mathrm{d}t\,,$ (13)
where we have used that
$(\boldsymbol{\omega}\mathrm{d}t\cdot\boldsymbol{J})^{2}=\sum_{ij}\omega_{i}\omega_{j}\mathrm{d}t^{2}J_{i}J_{j}\\ =\sum_{i}2D_{\text{R}}\mathrm{d}t(J_{i})^{2}=-4D_{\text{R}}\mathds{1}\mathrm{d}t\,.$ (14)
Equation 13 is the same as Eq. 9, which is seen by subtracting $\boldsymbol{u}(t)$ from both sides of Eq. 13 and dividing it by the differential $\mathrm{d}t$.
III.2 Discussion of the drift term
Before, we discuss the term $-\tau_{\text{R}}^{-1}\boldsymbol{u}(t)$ on the r.h.s. of Eq. 9, which appears to describe a drift. However, the term is normal to the constraining manifold and hence cannot generate a drift. Instead, it is needed to satisfy the constraint, i.e., to preserve the normalisation $|\boldsymbol{u}(t)|=1$ under the dynamics of Eq. 9. In order to prove this fact, we switch to the notation of stochastic (Itō) differentials and write $\mathrm{d}\boldsymbol{u}(t)=\dot{\boldsymbol{u}}(t)\mathrm{d}t$; the expression $\boldsymbol{\omega}(t)\mathrm{d}t$ is the differential of a scaled (and shifted  [^3] The argument given holds also in the presence of $\boldsymbol{T}_{\text{ext}}(t)$ in Eq. 6.) Wiener process. Then, Eq. 6 implies $\omega_{i}(t)\mathrm{d}t\,\omega_{j}(t)\mathrm{d}t=2D_{\text{R}}\delta_{ij}\mathrm{d}t$ for the Cartesian components $i$ and $j$ of $\boldsymbol{\omega}(t)$. Employing Itō’s formula, substituting $\mathrm{d}\boldsymbol{u}(t)$ by means of Eq. 9, and omitting the time arguments for brevity, one finds:
$\mathrm{d}(\boldsymbol{u}\cdot\boldsymbol{u})$ $=2\boldsymbol{u}\cdot\mathrm{d}\boldsymbol{u}+\mathrm{d}\boldsymbol{u}\cdot\mathrm{d}\boldsymbol{u}$
=(9)${{=}}-2\tau_{\text{R}}^{-1}\boldsymbol{u}\cdot\boldsymbol{u}\mathrm{d}t+(\boldsymbol{\omega}\mathrm{d}t\times\boldsymbol{u})\cdot(\boldsymbol{\omega}\mathrm{d}t\times\boldsymbol{u})$ )
$=-2\tau_{\text{R}}^{-1}|\boldsymbol{u}|^{2}\mathrm{d}t+\sum_{i}(\omega_{i}\mathrm{d}t)^{2}|\boldsymbol{u}|^{2}$
$\qquad-\sum_{ij}\omega_{i}\mathrm{d}t\,\omega_{j}\mathrm{d}t\,u_{i}\,u_{j}$
$=-2\tau_{\text{R}}^{-1}|\boldsymbol{u}|^{2}\mathrm{d}t+6D_{\text{R}}\mathrm{d}t|\boldsymbol{u}|^{2}-2D_{\text{R}}\mathrm{d}t\,|\boldsymbol{u}|^{2}$
$=2\left(2D_{\text{R}}-\tau_{\text{R}}^{-1}\right)|\boldsymbol{u}|^{2}\mathrm{d}t\,.$ (15)
Therefore, provided that $\tau_{\text{R}}^{-1}=2D_{\text{R}}$, it holds $\mathrm{d}|\boldsymbol{u}(t)|^{2}/\mathrm{d}t=0$ or, equivalently, $|\boldsymbol{u}(t)|=\mathrm{const}$.
III.3 Orientational autocorrelation function
The importance of the Itō term in Eq. 9 becomes clear again when computing the autocorrelation function of the orientation, $C_{1}(t):=\langle\boldsymbol{u}(t)\cdot\boldsymbol{u}(0)\rangle$, which is a simple exercise in the absence of an external torque $\boldsymbol{T}_{\text{ext}}(t)=0$. Multiplication of Eq. 9 by $\boldsymbol{u}(0)$, integrating over time, and averaging yields
$C_{1}(t)=\biggl{\langle}\int_{0}^{t}[\boldsymbol{\omega}(s)\times\boldsymbol{u}(s)]\cdot\boldsymbol{u}(0)\,\mathrm{d}s\biggr{\rangle}-\tau_{\text{R}}^{-1}\int_{0}^{t}C(s)\,\mathrm{d}s\,.$ (16)
Recalling that $\boldsymbol{\omega}(s)$ is a white noise process [Eq. 6] and independent of $\boldsymbol{u}(s)$ and $\boldsymbol{u}(0)$, the first term on the r.h.s. contains a properly formed Itō integral, which is zero on average. (The Stratonovich integral does not share this property.) Here, we have used that $\langle\omega(t)\rangle=\boldsymbol{T}_{\text{ext}}(t)=0$. The remaining integral equation has the expected solution
$C_{1}(t)=\exp(-t/\tau_{\text{R}})\,,$ (17)
which is seen, for example, by differentiating with respect to $t$ and solving the obtained ordinary differential equation for $C_{1}(t)$ with $C_{1}(0)=|\boldsymbol{u}(0)|^{2}=1$.
III.4 Equivalence of the Itō and Stratonovich forms
It remains to show the mathematical equivalence of Eqs. 7 and 9. Generally, given a diffusion process $\boldsymbol{X}(t)$ in $\mathbb{R}^{d}$, which is driven by a standard Wiener process $\boldsymbol{W}(t)$ in $\mathbb{R}^{n}$ scaled by an $d\times n$ matrix-valued coefficient function $\sigma(\boldsymbol{X}(t))$, the corresponding Stratonovich integral, denoted by $\hbox{}\circ\mathrm{d}\boldsymbol{W}(t)$ as usual, is related to the Itō integral via [79, 51]
$\int_{0}^{T}\sigma_{ij}(\boldsymbol{X}(t))\circ\mathrm{d}W_{j}(t)=\int_{0}^{T}\sigma_{ij}(\boldsymbol{X}(t))\,\mathrm{d}W_{j}(t)\\ +\frac{1}{2}\int_{0}^{T}[\partial_{k}\sigma(\boldsymbol{X}(t))]_{ij}\sigma_{kj}(\boldsymbol{X}(t))\,\mathrm{d}t\,,$ (18)
for some time $T>0$ and $\partial_{k}$ denoting the partial derivative w.r.t. to the component $X_{k}$. Omitting the integral signs in Eq. 18 (formally, taking a derivative w.r.t. $T$) yields the differential form of this relation in the common short-hand notation:
$\sigma_{ij}(\boldsymbol{X})\circ\mathrm{d}W_{j}=\sigma_{ij}(\boldsymbol{X})\,\mathrm{d}W_{j}+\frac{1}{2}[\partial_{k}\sigma(\boldsymbol{X})]_{ij}\sigma_{kj}(\boldsymbol{X})\,\mathrm{d}t\,.$ (19)
In order to apply it to Eq. 7, one identifies $\boldsymbol{X}(t)$ with $\boldsymbol{u}(t)$ and $\mathrm{d}\boldsymbol{W}(t)$ with $(2D_{\text{R}})^{-1/2}\boldsymbol{\omega}(t)\mathrm{d}t$. The noise strength $\sigma(\cdot)$ is $3\times 3$ matrix-valued, and it is convenient to use the above matrix form of the vector product: $\sigma(\boldsymbol{u})=-(2D_{\text{R}})^{1/2}\boldsymbol{u}\cdot\boldsymbol{J}$, which has the derivatives $\partial_{k}\sigma(\boldsymbol{u})=-(2D_{\text{R}})^{1/2}J_{k}$. It follows that
$[-\boldsymbol{u}(t)\cdot\boldsymbol{J}]_{ij}\circ\omega_{j}(t)\,\mathrm{d}t$ $=[-\boldsymbol{u}(t)\cdot\boldsymbol{J}]_{ij}\,\omega_{j}(t)\,\mathrm{d}t$
$\qquad+\frac{2D_{\text{R}}}{2}(-J_{k})_{ij}[-\boldsymbol{u}(t)\cdot\boldsymbol{J}]_{kj}\,\mathrm{d}t$
$=\boldsymbol{\omega}(t)\,\mathrm{d}t\times\boldsymbol{u}(t)-2D_{\text{R}}\boldsymbol{u}(t)\,\mathrm{d}t\,;$ (20)
in the last step, we have simplified the last term as follows:
$(-J_{k})_{ij}[-\boldsymbol{u}(t)\cdot\boldsymbol{J}]_{kj}=(J_{k})_{ij}(J_{l})_{kj}u_{l}(t)\\ =\varepsilon_{kij}\varepsilon_{lkj}u_{l}(t)=(\delta_{kl}\delta_{ik}-\delta_{kk}\delta_{il})u_{l}(t)=-2u_{i}(t)\,;$ (21)
$\delta_{ij}$ is the Kronecker symbol. After division of Eq. 20 by $\mathrm{d}t$, it is clear that, for $\tau_{\text{R}}^{-1}=2D_{\text{R}}$, the r.h.s. of the Stratonovich SDE (7) corresponds to the r.h.s. of the Itō SDE (9) as claimed.
III.5 Two-dimensional case
The geometric derivation of Eq. 9 has a straightforward generalisation to $d=2$ dimensions. We note that the angular velocity $\boldsymbol{\omega}(t)$, being an axial vector, has $n:=d(d-1)/2$ components since it can be represented as a skew-symmetric $d\times d$ matrix. In particular for $d=2$, it has only a single component $\omega_{z}$ (which may be chosen as the $z$-component in a three-dimensional embedding of the motion). Correspondingly, the Lie algebra $\mathfrak{so}(2)$ is generated by a single matrix, e.g., $J_{z}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$. With this, the cross product in Eq. 12 is to be understood as
$\boldsymbol{\omega}\times\boldsymbol{v}=(\boldsymbol{\omega}\cdot\boldsymbol{J})\boldsymbol{v}=\omega_{z}J_{z}\boldsymbol{v}$ (22)
for two-dimensional vectors $\boldsymbol{v}$. The reasoning from Eqs. 10 to 13 remains unchanged except for the last step, where one has to replace $-\sum_{i}(J_{i})^{2}=2\cdot\mathds{1}$ by $-(J_{z})^{2}=\mathds{1}$. Hence, Eq. 9 applies also in $d=2$ dimensions, but with $\tau_{\text{R}}^{-1}=D_{\text{R}}$.
For the proof of the equivalence of the Itō and Stratonovich forms (Section III.4), the following changes apply: The noise strength $\sigma(\boldsymbol{u})=-(2D_{\text{R}})^{1/2}J_{z}\boldsymbol{u}$ is a $2\times 1$ matrix with derivative $\partial_{k}\sigma(\boldsymbol{u})=-(2D_{\text{R}})^{1/2}J_{z}\boldsymbol{e}_{k}$; here, $\boldsymbol{e}_{k}$ are the Cartesian unit vectors, $(\boldsymbol{e}_{k})_{j}=\delta_{kj}$. Second, the calculation in Eq. 21 is replaced by
$(2D_{\text{R}})^{-1}[\partial_{k}\sigma(\boldsymbol{u})]_{ij}\sigma(\boldsymbol{u})_{kj}=(-J_{z}\boldsymbol{e}_{k})_{ij}[-J_{z}\boldsymbol{u}(t)]_{kj}\\ =(J_{z})_{ik}(J_{z})_{kl}u_{l}(t)=[(J_{z})^{2}\boldsymbol{u}(t)]_{i}=-u_{i}(t)\,,$ (23)
where the sums run over rows $k,l\in\{1,2\}$ and column $j=1$.
IV Geometric numerical integration scheme
IV.1 Algorithm
Above, we have shown that Eq. 10 is equivalent to Eq. 9. The latter is a standard Itō stochastic differential equation and may be integrated numerically with the Euler–Maruyama scheme or some higher-order scheme [81, 51], introducing some finite integration time step $\Delta t$. However, these schemes would satisfy the constraint $|\boldsymbol{u}(t)|=1$ only asymptotically, for $\Delta t\to 0$.
Alternatively, Eq. 10 suggests to implement a geometric integration scheme for Brownian motion on the unit sphere as a sequence of rotations with finite random angles, which reads for a single integration step:
$\boldsymbol{u}(t+\Delta t)=\text{e}^{\Delta\boldsymbol{\Omega}(t)\cdot\boldsymbol{J}}\boldsymbol{u}(t)$ (24)
for yet to be determined random vectors $\Delta\boldsymbol{\Omega}(t)$, recalling that $\text{e}^{\vartheta\boldsymbol{n}\cdot\boldsymbol{J}}$ is a rotation matrix with axis $\boldsymbol{n}:=\Delta\boldsymbol{\Omega}(t)/|\Delta\boldsymbol{\Omega}(t)|$ and angle $\vartheta:=|\Delta\boldsymbol{\Omega}(t)|$. The explicit action of this rotation on the unit vector $\boldsymbol{u}=\boldsymbol{u}(t)$ is given for $d=3$ by Rodrigues formula,
$\boldsymbol{u}(t+\Delta t)=\cos(\vartheta)\boldsymbol{u}-\sin(\vartheta)\,\boldsymbol{u}\times\boldsymbol{n}+[1-\cos(\vartheta)](\boldsymbol{u}\cdot\boldsymbol{n})\boldsymbol{n}\,.$ (25)
We note that, nowadays, the trigonometric functions are evaluated by dedicated special function units on typical computing hardware, either CPUs or GPUs, and the appearance of such functions in the algorithm is not a performance issue anymore.
It remains to specify the statistics of $\Delta\boldsymbol{\Omega}(t)$. Comparison of Eq. 24 and Eq. 10 yields for small time steps that
$\Delta\boldsymbol{\Omega}(t)\simeq\boldsymbol{\omega}(t)\Delta t\,;\quad\Delta t\to 0.$ (26)
With this choice, time integration of Eq. 6 suggests that $\Delta\boldsymbol{\Omega}(t)$ is a Gaussian vector with independent components. Clearly, the rotation of $\boldsymbol{u}$ about its own axis has no effect, which also follows from setting $\boldsymbol{n}=\boldsymbol{u}$ in Eq. 25. Thus, it is sufficient to choose $\Delta\boldsymbol{\Omega}$ in the plane perpendicular to $\boldsymbol{u}$:
$\Delta\boldsymbol{\Omega}=\Omega_{1}\boldsymbol{e}_{1}+\Omega_{2}\boldsymbol{e}_{2}$ (27)
for unit vectors $(\boldsymbol{e}_{1},\boldsymbol{e}_{2},\boldsymbol{u})$ forming a trihedron. The random coefficients $\Omega_{1}$ and $\Omega_{2}$ are independent Gaussian variables with means and variances given by
$\langle\Omega_{i}\rangle=\boldsymbol{e}_{i}\cdot\zeta_{R}^{-1}\boldsymbol{T}_{\text{ext}}(t)\quad\text{and}\quad\operatorname{Var}[\Omega_{i}]=2D_{\text{R}}\Delta t$ (28)
for $i=1,2$, also including a possible external torque.
Inputs: $\boldsymbol{u}=\boldsymbol{u}(t)$, $\boldsymbol{T}_{\text{ext}}(t)$, $D_{\text{R}}=k_{\text{B}}T\zeta_{R}^{-1}$, $\Delta t$
Output: $\boldsymbol{u}(t+\Delta t)$
(i) construct the vector $\boldsymbol{e}_{1}\perp\boldsymbol{u}$ and let $\boldsymbol{e}_{2}=\boldsymbol{u}\times\boldsymbol{e}_{1}$ (ii) draw normally distributed random coefficients $\Omega_{1}$ and $\Omega_{2}$ such that $\langle\Omega_{i}\rangle=\boldsymbol{e}_{i}\cdot\zeta_{R}^{-1}\boldsymbol{T}_{\text{ext}}(t)$ and $\operatorname{Var}[\Omega_{i}]=2D_{\text{R}}\Delta t$ (iii) compute $\boldsymbol{n}=\Delta\boldsymbol{\Omega}/|\Delta\boldsymbol{\Omega}|$ and $\vartheta=|\Delta\boldsymbol{\Omega}|$ and evaluate $\sin(\vartheta)$ and $\cos(\vartheta)$ (iv) obtain $\boldsymbol{u}(t+\Delta t)=\cos(\vartheta)\boldsymbol{u}-\sin(\vartheta)\,\boldsymbol{u}\times\boldsymbol{n}$
Table 1: Algorithm for the geometric integration scheme [Eq. 24] using Gaussian distributed random rotations [Eq. 28].
In summary, each time step of the geometric integration scheme [Eqs. 24, 25, 26 and 28] consists of the algorithm listed in Table 1. In step (iv), we have simplified Eq. 25, noting that $\boldsymbol{u}\perp\boldsymbol{n}$. Optionally, after step (iv), one may normalise the unit vector $\boldsymbol{u}(t+\Delta t)$ to avoid a possible drift of $|\boldsymbol{u}|$ after a large number of integration steps, which can result from round-off errors due to limited floating-point precision. We emphasise again that the time discretisation in this scheme preserves the normalisation exactly.
IV.2 Estimation of the numerical error
The form of Eq. 24 admits an exact representation of the solution of the SDE (10) whereas Eq. 26 describes merely an approximation because
$\exp(\Delta\boldsymbol{\Omega}_{A}\cdot\boldsymbol{J})\exp(\Delta\boldsymbol{\Omega}_{B}\cdot\boldsymbol{J})\neq\exp((\Delta\boldsymbol{\Omega}_{A}+\Delta\boldsymbol{\Omega}_{B})\cdot\boldsymbol{J})$ (29)
for $\Delta\boldsymbol{\Omega}_{A}$ and $\Delta\boldsymbol{\Omega}_{B}$ of subsequent integration steps. In order to estimate the discretisation error, we compare the obtained distribution of $\boldsymbol{u}(t+\Delta t)$ with the propagator of the dynamics of $\boldsymbol{u}(t)$, putting $\boldsymbol{T}_{\text{ext}}(t)=0$ for simplicity. Then, the free propagator is known from the solution of the corresponding Fokker–Planck equation and, assuming that $\boldsymbol{u}(0)$ is sampled in equilibrium, it reads in terms of the correlation functions [47]
$C_{\ell}(t)$ $:=\langle P_{\ell}\boldsymbol{(}\boldsymbol{u}(t)\cdot\boldsymbol{u}(0)\boldsymbol{)}\rangle$
$=\text{e}^{-\ell(\ell+1)D_{\text{R}}\Delta t}\,,\qquad\ell=0,1,2,\dots\,,$ (30)
where $P_{\ell}(\cdot)$ denotes the Legendre polynomial of degree $\ell$. Since $\boldsymbol{u}(t)$ is a Markov process, it is sufficient to consider the first step from $t=0$ to $t=\Delta t$. On the other hand, the numerical scheme [Eq. 25] yields $\boldsymbol{u}(t)\cdot\boldsymbol{u}(0)=\cos(\vartheta)$ and thus
$\widehat{C}_{\ell}(t)=\langle P_{\ell}\boldsymbol{(}\cos(|\Delta\boldsymbol{\Omega}|)\boldsymbol{)}\rangle_{\Delta\boldsymbol{\Omega}}\,,$ (31)
where we have used that $\boldsymbol{n}$ is orthogonal to $\boldsymbol{u}(0)$ and $|\boldsymbol{u}(0)|=1$ and the average is taken with respect to the distribution of $\Delta\boldsymbol{\Omega}$ given in Eq. 28. Inserting the Gaussian distributions for $\Omega_{1}$ and $\Omega_{2}$ and transforming to the angle $\vartheta=(\Omega_{1}^{2}+\Omega_{2}^{2})^{1/2}$, one obtains
$\widehat{C}_{\ell}(t)=\int_{0}^{\infty}P_{\ell}\boldsymbol{(}\cos(\vartheta)\boldsymbol{)}\,\frac{1}{2D_{\text{R}}t}\,\text{e}^{-\vartheta^{2}/4D_{\text{R}}t}\,\vartheta\mathrm{d}\vartheta\,.$ (32)
This integral can be expressed in terms of Dawson’s $F$-function [85]. For our purposes, however, it suffices to expand the integrand in $\vartheta=0$ since the distribution is sharply peaked for small $D_{\text{R}}t\ll 1$:
$\widehat{C}_{\ell}(t)$ $=\int_{0}^{\infty}\left[1-\frac{\ell(\ell+1)}{4}\vartheta^{2}+\frac{1}{24}a_{2}(\ell)\vartheta^{4}+O(\vartheta^{6})\right]\,$
×$\qquad\times\frac{1}{4D_{\text{R}}t}\,\text{e}^{-\vartheta^{2}/4D_{\text{R}}t}\,\mathrm{d}(\vartheta^{2})$
$=1-\ell(\ell+1)D_{\text{R}}t+\frac{4}{3}a_{2}(\ell)(D_{\text{R}}t)^{2}+O\left((D_{\text{R}}t)^{4}\right)$ (33)
where $\alpha_{2}(\ell)=\ell(\ell+1)[3\ell(\ell+1)-2]/8$. From the comparison with Eq. 30 and writing $\Delta t$ again instead of $t$, one finds the relative numerical error of the correlation functions $C_{\ell}(t)$ after one integration step of length $\Delta t$:
$\frac{\widehat{C}_{\ell}(\Delta t)-C_{\ell}(\Delta t)}{C_{\ell}(\Delta t)}=-\frac{\ell(\ell+1)}{3}\,(D_{\text{R}}\Delta t)^{2}+O\left((D_{\text{R}}\Delta t)^{3}\right)\,,$ (34)
asymptotically as $D_{\text{R}}\Delta t\to 0$. In order to integrate $\boldsymbol{u}(0)$ up to time $t=N\Delta t$ one applies the one-step propagator [Eq. 24] successively $N$ times. Correspondingly, the relative discretisation error [Eq. 34] adds up $N$ times so that the overall error obeys
$|\widehat{C}_{\ell}(t)-C_{\ell}(t)|\simeq\frac{D_{\text{R}}\Delta t}{3}\,\ell(\ell+1)D_{\text{R}}t\,\text{e}^{-\ell(\ell+1)D_{\text{R}}t}$ (35)
upon $D_{\text{R}}\Delta t\to 0$, which we have verified in numerical tests. Moreover, the $t$-dependent factor is bounded and we can estimate the error for arbitrary times:
$\sup_{t\in[0,\infty)}|\widehat{C}_{\ell}(t)-C_{\ell}(t)|=\frac{D_{\text{R}}\Delta t}{3\text{e}}+O\left((D_{\text{R}}\Delta t)^{2}\right)\,.$ (36)
One concludes that the geometric integrator shows a globally weak convergence of order 1 in the integration time step $\Delta t$.
IV.3 A weakly exact integration scheme
A weakly exact integration scheme (i.e., one that yields the correct statistics of the solution) is obtained upon replacing Eq. 28 by the true statistics of $\Delta\boldsymbol{\Omega}$, which tends to a Gaussian only for $D_{\text{R}}\Delta t\ll 1$. In particular, a series representation of the distribution of the rotation angle $\vartheta=|\Delta\boldsymbol{\Omega}|$, restricted to $0\leqslant\vartheta\leqslant\pi$, follows from its angular moments [Eq. 30]:
$p(\vartheta,t)=\sum_{\ell\geqslant 0}\left(\ell+\tfrac{1}{2}\right)C_{\ell}(t)P_{\ell}\boldsymbol{(}\cos(\vartheta)\boldsymbol{)}\sin(\vartheta)\,.$ (37)
With this, $C_{\ell}(t)=\int_{0}^{\pi}P_{\ell}\boldsymbol{(}\cos(\vartheta)\boldsymbol{)}\,p(\vartheta,t)\,\mathrm{d}\vartheta$ and reweighting the angle $\vartheta$ in step (iii) of the algorithm in Table 1 such that it samples $p(\vartheta)$ yields numerical correlation functions $\widehat{C}_{\ell}(t)$ [Eq. 31] which are equal to the exact solution $C_{\ell}(t)$ for all $t$. A truncation of the series for $p(\vartheta,t)$ by keeping only terms with $\ell\leqslant\ell_{\text{max}}$ yields $\widehat{C}_{\ell}(t)=C_{\ell}(t)$ for $\ell\leqslant\ell_{\text{max}}$ and $\widehat{C}_{\ell}(t)=0$ otherwise. We note that $p(\vartheta,t)$ is similar in shape, but different from the van-Mises-Fisher distribution, $p_{\text{vMF}}(\vartheta,t)\propto\exp(\cos(\vartheta)/2D_{\text{R}}t)\sin(\vartheta)$, which decays too slowly as $\vartheta\to\pi$ and would yield a larger integration error than given in Eq. 34.
Numerically, the direct sampling from the (truncated or untruncated) distribution (37) may be implemented via inverse transform sampling: $F^{-1}(Z)$ samples the density $p(\vartheta,\Delta t)$ if $Z$ is a uniformly distributed random variable on the interval $[0,1]$ and the function $F^{-1}(\cdot)$ is the (numerically pre-computed) inverse of the cumulative distribution function of $\vartheta$, i.e., $F(\vartheta)=\int_{0}^{\vartheta}p(\vartheta^{\prime},\Delta t)\,\mathrm{d}\vartheta^{\prime}$ for $\vartheta\in[0,\pi]$. Alternatively, one can imagine to start from a Gaussian distributed variable $Z$ and to use $Z=|\Delta\boldsymbol{\Omega}|$ from the original step (iii) of the algorithm. An efficient implementation of this exact variant of the integration scheme is left for future research.
IV.4 Comparison to the ad hoc scheme
Figure 1: Numerically obtained correlation functions $\widehat{C}_{\ell}(t)$ for $\ell=1$ (circles) and $\ell=2$ (diamonds) using the geometric integrator (algorithm in Table 1, closed symbols) and the ad hoc scheme [Eqs. (38), open symbols], which is based on the Euler–Maruyama discretisation of the Langevin equation (7) with an a posteriori normalisation of the orientation vector. Colours distinguish different integration time steps $\Delta t$, numerical data are available only for correlation times $t$ that are a multiple of $\Delta t$, and coloured lines serve as guides to the eye. The solid and dashed black lines show the exact solutions $C_{\ell}(t)$ for $\ell=1$ and $\ell=2$, respectively [Eq. 30].
For completeness, we also give a brief analysis of the numerical scheme suggested in [66], which combines an Euler–Maruyama step for the Stratonovich SDE (7) with a post-normalisation step of the unit vector; the second step corresponds to an orthogonal projection onto the constraining manifold, $|\boldsymbol{u}(t+\Delta t)|=1$. This scheme corresponds to the geometric algorithm in Table 1 with step (iv) replaced by:
$\boldsymbol{u}^{\prime}$ $=\boldsymbol{u}(t)+\Delta\boldsymbol{\Omega}(t)\times\boldsymbol{u}(t)\,,$ (38a)
$\boldsymbol{u}(t+\Delta t)$ $=\boldsymbol{u}^{\prime}/|\boldsymbol{u}^{\prime}|\,;$ (38b)
as before, $\Delta\boldsymbol{\Omega}(t)$ is a Gaussian random vector satisfying Eq. 28. The first equation may also be obtained from Eq. 25 by expanding in $\vartheta=0$ and keeping only terms $O(\vartheta)$.
A calculation similar to the one in Section III.2 shows that the post-normalisation in Eq. 38b is equivalent (in distribution) to adding the Itō drift term to Eq. 38a up to linear order in $\Delta t\to 0$:
$\boldsymbol{u}(t+\Delta t)$ $=\frac{\boldsymbol{u}(t)+\Delta\boldsymbol{\Omega}(t)\times\boldsymbol{u}(t)}{(1+|\Delta\boldsymbol{\Omega}(t)\times\boldsymbol{u}(t)|^{2})^{1/2}}$
$\simeq\boldsymbol{u}(t)+\Delta\boldsymbol{\Omega}(t)\times\boldsymbol{u}(t)-2D_{\text{R}}\boldsymbol{u}(t)\Delta t\,.$ (39)
We conclude that the ad hoc scheme has the same convergence properties as the Euler–Maruyama scheme for the Itō–Langevin equation [Eq. 9]; in particular, it also exhibits weak convergence of order 1. However, the magnitude of the error is significantly larger compared to the geometric scheme (Table 1). This is demonstrated by numerical results for the correlation functions $\widehat{C}_{1}(t)$ and $\widehat{C}_{2}(t)$ obtained from both schemes (Fig. 1). For an integration time step of $\Delta t=0.1D_{\text{R}}^{-1}$, the results from the ad hoc scheme deviate already considerably from the exact solutions, whereas merely insignificant deviations are visible for $\Delta t=0.3D_{\text{R}}^{-1}$ using the geometric scheme.
Figure 2: Probability distribution $p(\vartheta,\Delta t)$ of the angle $\vartheta$ between the initial and final orientations after an integration time step of length $\Delta t=0.3D_{\text{R}}^{-1}$ for the geometric scheme with Gaussian rotational vectors [Eq. 40, green line] and for the ad hoc scheme, combining an Euler–Maruyama step with an orthogonal projection [Eq. 43, red line]. The grey line shows the exact propagator of rotational Brownian motion [Eq. 37].
For the geometric scheme with Gaussian rotation vectors, the wrapped rotation angle $0\leqslant\vartheta\leqslant\pi$ is distributed with the probability density [see Eq. 32]:
$p_{\text{Gauss}}(\vartheta,\Delta t)$ $=\sum_{k\geqslant 0}\frac{\vartheta+2\pi k}{2D_{\text{R}}\Delta t}\,\text{e}^{-(\vartheta+2\pi k)^{2}/4D_{\text{R}}\Delta t}\,.$ (40)
For most practical purposes, it is sufficient to restrict the sum to the $k=0$ term.
In the case of the ad hoc scheme, the final rotation angle is given via [see Eq. 39]:
$\cos(\vartheta)$ $=\boldsymbol{u}(t+\Delta t)\cdot\boldsymbol{u}(t)$
$=\bigl{(}1+|\Delta\boldsymbol{\Omega}(t)|^{2}\bigr{)}^{-1/2}\,,$ (41)
making use of $|\boldsymbol{u}(t)|=1$ and $\Delta\boldsymbol{\Omega}(t)\perp\boldsymbol{u}(t)$. One finds for the distribution of $z=\cos(\vartheta)$:
$\bar{p}_{\text{ad hoc}}(z,\Delta t)$ $=\int_{0}^{\infty}\delta\left(z-\frac{1}{\sqrt{1+\vartheta^{2}}}\right)\,\frac{\text{e}^{-\vartheta^{2}/4D_{\text{R}}\Delta t}}{2D_{\text{R}}\Delta t}\,\vartheta\mathrm{d}\vartheta$
$=\frac{z^{-3}}{2D_{\text{R}}\Delta t}\text{e}^{-(z^{-2}-1)/4D_{\text{R}}\Delta t}\,,\quad z>0.$ (42)
Zero or negative values, $z\leqslant 0$, cannot be attained in this scheme since $\boldsymbol{u}^{\prime}$ is in the tangent plane of $\boldsymbol{u}$ and thus its projection cannot reach the lower half-sphere. It follows that the probability density of the angle $20\leqslant\vartheta<\pi/$ 2 reads:
$p_{\text{ad hoc}}(\vartheta,\Delta t)=\frac{\sin(\vartheta)}{2D_{\text{R}}\Delta t\cos(\vartheta)^{3}}\,\exp\left(-\frac{\cos(\vartheta)^{-2}-1}{4D_{\text{R}}\Delta t}\right)\,.$ (43)
One observes that, for small rotation angles,
$p_{\text{ad hoc}}(\vartheta\to 0,\Delta t)=p_{\text{Gauss}}(\vartheta,\Delta t)\,\bigl{[}1+O(\vartheta^{2}/D_{\text{R}}\Delta t)\bigr{]}\,.$ (44)
The obtained distributions $p(\vartheta,\Delta t)$ are the propagators of the various integration schemes investigated. Their graphical comparison in Fig. 2 for $\Delta t=0.3D_{\text{R}}^{-1}$ shows that the geometric scheme with Gaussian rotation vectors (Table 1) closely follows the exact distribution [Eq. 37], despite the comparably large time step. In contrast, the propagator of the ad hoc scheme deviates significantly, which underpins the observations made before for the correlation functions (Fig. 1).
V Extension to axisymmetric particles
For the extension to axisymmetric particles, let $\boldsymbol{u}(t)$ denote the symmetry axis under rotations. Then, the friction tensor and moment of inertia in the spaced-fixed frame attain the forms [71]:
$\boldsymbol{\zeta}_{R}(t)$ $=\zeta_{R}^{\parallel}\mathcal{P}_{\boldsymbol{u}}(t)+\zeta_{R}^{\perp}\mathcal{Q}_{\boldsymbol{u}}(t)\,,$ (45)
$\boldsymbol{I}(t)$ $=I_{\parallel}\mathcal{P}_{\boldsymbol{u}}(t)+I_{\perp}\mathcal{Q}_{\boldsymbol{u}}(t)\,,$ (46)
where $\mathcal{P}_{\boldsymbol{u}}(t)=\boldsymbol{u}(t)\otimes\boldsymbol{u}(t)$ is the orthogonal projector on the instantaneous direction $\boldsymbol{u}(t)$ and $\mathcal{Q}_{\boldsymbol{u}}(t)=\mathds{1}-\mathcal{P}_{\boldsymbol{u}}(t)$ its complement. For Brownian particles, the argument of time scale separation leading to Eq. 4 applies analogously since the matrix norms of $\boldsymbol{\zeta}_{R}(t)$ and $\boldsymbol{I}(t)$ remain unchanged under the rotational transformations. Thus, in the overdamped limit, Eq. 5 holds with a time-dependent, stochastic friction tensor $\boldsymbol{\zeta}_{R}(t)$. Multiplying the equation with the reciprocal tensor,
$\boldsymbol{\zeta}_{R}(t)^{-1}=(\zeta_{R}^{\parallel})^{-1}\mathcal{P}_{\boldsymbol{u}}(t)+(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t)\,,\$ (47)
yields
$\boldsymbol{\omega}(t)=\boldsymbol{\zeta}_{R}(t)^{-1}\boldsymbol{T}_{\text{ext}}(t)+\boldsymbol{\zeta}_{R}(t)^{-1}\boldsymbol{\xi}(t)\,.$ (48)
For application to the orientation vector, we note that only the component perpendicular to the orientation, $\boldsymbol{\omega}_{\perp}(t)=\mathcal{Q}_{\boldsymbol{u}}(t)\,\boldsymbol{\omega}(t)$, is relevant for the cross product in Eq. 9; it reads
$\boldsymbol{\omega}_{\perp}(t)=(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t)\boldsymbol{T}_{\text{ext}}(t)+(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t)\boldsymbol{\xi}(t)\,,$ (49)
which contains only the friction coefficient $\zeta_{R}^{\perp}$. Also, the evolution of $\boldsymbol{\omega}_{\perp}(t)$ does not depend on the remaining component $\boldsymbol{\omega}_{\parallel}(t)=\mathcal{P}_{\boldsymbol{u}}(t)\,\boldsymbol{\omega}(t)$, parallel to the particle axis. This allows one to replace $\boldsymbol{\omega}(t)$ in Eq. 9 by a process $\tilde{\boldsymbol{\omega}}(t)$, which satisfies Eq. 5 for isotropic friction, i.e., to set $\zeta_{R}^{\parallel}=\zeta_{R}^{\perp}=\zeta_{R}$. As an important consequence, the above discussion and analysis of the overdamped dynamics of $\boldsymbol{u}(t)$ for spherical particles carries over unchanged to axisymmetric particles. In particular, the orientation vector performs a homogeneous Brownian motion on the sphere (for constant $\boldsymbol{T}_{\text{ext}}$) and the described numerical integration scheme can be employed without modifications upon choosing $\zeta_{R}=\zeta_{R}^{\perp}$.
The last point may also be seen from a different angle: since the white noise $\boldsymbol{\xi}(t^{\prime})$ and the orientation $\boldsymbol{u}(t)$ are independent for $t^{\prime}\geqslant t$, the term $\mathcal{Q}_{\boldsymbol{u}}(t)\boldsymbol{\xi}(t)$ in Eq. 49 is equivalent (in probability) to an effectively two-dimensional, Gaussian white noise process that, for each time $t$, is rotated into the plane perpendicular to $\boldsymbol{u}(t)$. This construction is already reflected in the algorithm when choosing the rotation angle in Eq. 27.
For completeness, we specify also the mean and covariance of $\boldsymbol{\omega}_{\perp}(t)$ for $t^{\prime}\geqslant t$, conditioned on the value of $\boldsymbol{u}(t)$ [cf. Eq. 6]:
$\langle\boldsymbol{\omega}_{\perp}(t)|\boldsymbol{u}(t)\rangle$ $=(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t)\boldsymbol{T}_{\text{ext}}(t)\,,$ (50a)
$\langle\delta\boldsymbol{\omega}_{\perp}(t^{\prime})\otimes\delta\boldsymbol{\omega}_{\perp}(t)|\boldsymbol{u}(t)\rangle$ $=2D_{R}^{\perp}\mathcal{Q}_{\boldsymbol{u}}(t)\,\delta(t^{\prime}-t)\,,$ (50b)
where the condition is denoted on the right of the bar and $\delta\boldsymbol{\omega}_{\perp}(t):=\boldsymbol{\omega}_{\perp}(t)-\langle\boldsymbol{\omega}_{\perp}(t)|\boldsymbol{u}(t)\rangle$. In the second equation, the perpendicular part of the $\boldsymbol{u}(t)$-dependent rotational diffusion tensor enters:
$\boldsymbol{D}_{R}(t):=D_{R}^{\parallel}\mathcal{P}_{\boldsymbol{u}}(t)+D_{R}^{\perp}\mathcal{Q}_{\boldsymbol{u}}(t)\,,\$ (51)
with constants $D_{R}^{\parallel,\perp}=k_{\text{B}}T/\zeta_{R}^{\parallel,\perp}$. A derivation and further explanations of Eqs. 50a and 50b can be found in Appendix A.
VI Summary and conclusions
Brownian motion on the unit sphere is a central ingredient to models of axisymmetric colloidal particles, in which it describes the stochastic evolution of the symmetry axis; examples are oblate and prolate ellipsoids, but also active Brownian particles, e.g., Janus spheres, where such an axis is given by the propulsion direction. Here, we have constructed the corresponding overdamped stochastic process as a sequence of infinitesimal random rotations [Eq. 10]. This construction is known as McKean–Gangolli injection [83], which transforms a stochastic process on the Lie algebra — here, $\mathfrak{so}(3)$ — into the generated Lie group — here, the rotation group $SO(3)$. Within Itō calculus, it is then straightforward to derive a stochastic differential (i.e., Langevin) equation of the process [Eq. 9]. Transforming the equation to its equivalent form in Stratonovich calculus [Eq. 7] yields the same equation as is obtained heuristically from the underdamped Langevin equation of rotational Brownian motion [Eqs. 2 and 4]. In particular, the latter form is free from the apparent, noise-induced drift term present in the Itō form. We have shown that, indeed, this additional term does not generate a drift (bias) of the rotational motion; but rather, it is needed to preserve the normalisation of the orientation vector, $|\boldsymbol{u}(t)|=1$, under the stochastic dynamics.
In generic numerical integration schemes of the overdamped Langevin equation (in either form), such as the Euler–Maruyama and Milstein schemes, the normalisation constraint is satisfied only asymptotically for vanishing time step, ${\Delta t\to 0}$, depending on their strong order of convergence. As an alternative, we have proposed a geometric integration scheme [Eq. 24], which is based on finite random rotations and which satisfies the constraint exactly for any time step $\Delta t$. Combining this approach with Gaussian distributed random rotation angles yields an immediate and simple implementation of this geometric integrator (Table 1), which we have shown to converge weakly at order 1, i.e., $O(\Delta t)$, and which can also be combined with a deterministic bias, e.g., due to an external torque or to model circle swimmers. Moreover, we have outlined an improved algorithm that exactly generates the (free) propagator of rotational Brownian motion for an arbitrary time step $\Delta t$ if an advanced sampling of the rotation angle is used [see Eq. 37]. Finally, we have demonstrated numerically and by deriving the corresponding propagators that the proposed Gaussian geometric scheme converges more rapidly than a widely used ad hoc scheme [Eqs. (38)]; the smaller prefactor of the discretisation error allows for an integration time step that is an order of magnitude larger (Fig. 1).
Appendix A Statistics of the projected angular velocity
For a derivation of the first two moments of the projected angular velocity $\omega_{\perp}(t)$, given in Eqs. 50a and 50b, we switch to the integrated noise process,
$\boldsymbol{\Omega}_{\perp}(t)=\int_{0}^{t}\boldsymbol{\omega}_{\perp}(t^{\prime})\,\mathrm{d}t^{\prime}\,,$ (52)
which in differential form reads $\mathrm{d}\boldsymbol{\Omega}_{\perp}(t)=\boldsymbol{\omega}_{\perp}(t)\,\mathrm{d}t$. It is a proper Itō processes and, using Eqs. 3 and 49, it obeys the following (Itō) SDE in standard form [51, 52]:
$\mathrm{d}\boldsymbol{\Omega}_{\perp}(t)=(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t)\boldsymbol{T}_{\text{ext}}(t)\,\mathrm{d}t+(2D_{R}^{\perp})^{1/2}\mathcal{Q}_{\boldsymbol{u}}(t)\,\mathrm{d}\boldsymbol{W}(t)\,,$ (53)
which is driven by a standard three-dimensional Wiener process $\boldsymbol{W}(t)$ normalised such that $\langle|\boldsymbol{W}(t)|^{2}\rangle=3t$.
The mean value of a small increment $\Delta\boldsymbol{\Omega}_{\perp}(t):=\boldsymbol{\Omega}_{\perp}({t+\Delta t})-\boldsymbol{\Omega}_{\perp}(t)$, conditioned on $\boldsymbol{u}(t)$, is given by the drift part only:
$\mathrlap{E[\Delta\boldsymbol{\Omega}_{\perp}(t)|\boldsymbol{u}(t)]}\hskip 30.00005pt$
$=E\biggl{[}\int_{t}^{t+\Delta t}(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t^{\prime})\boldsymbol{T}_{\text{ext}}(t^{\prime})\,\mathrm{d}t^{\prime}\biggl{|}\boldsymbol{u}(t)\biggr{]}$
$\simeq(\zeta_{R}^{\perp})^{-1}\mathcal{Q}_{\boldsymbol{u}}(t)\boldsymbol{T}_{\text{ext}}(t)\Delta t\,;\quad\Delta t\to 0.$ (54)
Here, $E[\,\cdot\,|\boldsymbol{u}(t)]$ denotes the conditional expectation [52] with respect to the values of $\boldsymbol{u}(t)$ and, to obtain the second line, we have used continuity of the integrand. The noise part of $\mathrm{d}\boldsymbol{\Omega}_{\perp}(t)$ does not contribute since $\Delta\boldsymbol{W}(t):=\boldsymbol{W}(t+\Delta t)-\boldsymbol{W}(t)$ for $\Delta t>0$ is independent of $\boldsymbol{u}(t)$ and thus:
$E\biggl{[}\int_{t}^{t^{\prime}+\Delta t}\mathcal{Q}_{\boldsymbol{u}}(t^{\prime})\,\mathrm{d}\boldsymbol{W}(t^{\prime})\biggl{|}\boldsymbol{u}(t)\biggr{]}=0.$ (55)
Switching to the derivative of $\boldsymbol{\Omega}_{\perp}(t)$ in Eq. 54 yields Eq. 50a.
For the covariance function of $\boldsymbol{\Omega}_{\perp}(t)$, we compensate the drift and consider the rescaled process $\mathrm{d}\boldsymbol{W}^{\perp}(t)=\mathcal{Q}_{\boldsymbol{u}}(t)\,\mathrm{d}\boldsymbol{W}(t)$ such that $\int_{0}^{t}\delta\boldsymbol{\omega}_{\perp}(t^{\prime})\,\mathrm{d}t^{\prime}=(2D_{R}^{\perp})^{1/2}\boldsymbol{W}^{\perp}(t)$. Non-overlapping increments of $\boldsymbol{W}^{\perp}(t)$ are uncorrelated, which is seen by choosing $\Delta t^{\prime},\Delta t>0$ and $t^{\prime}>t+\Delta t$ and performing basic manipulations of conditional expectation:
$\mathrlap{E[\Delta W^{\perp}_{i}(t^{\prime})\Delta W^{\perp}_{j}(t)|\boldsymbol{u}(t)]}\hskip 30.00005pt$
$=E[E[\Delta W^{\perp}_{i}(t^{\prime})|\boldsymbol{u}(t^{\prime})]\,\Delta W^{\perp}_{j}(t)|\boldsymbol{u}(t)]$
$=0\,.$ (56)
Choosing the same time interval for both increments, it holds:
$E[\Delta W^{\perp}_{i}(t)\,\Delta W^{\perp}_{j}(t)|\boldsymbol{u}(t)]$ $=E\biggl{[}\int_{t}^{t+\Delta t}[\mathcal{Q}_{\boldsymbol{u}}(t^{\prime})]_{ik}\mathrm{d}W_{k}(t^{\prime})\int_{t}^{t+\Delta t}[\mathcal{Q}_{\boldsymbol{u}}(t^{\prime\prime})]_{jl}\mathrm{d}W_{l}(t^{\prime\prime})\biggl{|}\boldsymbol{u}(t)\biggr{]}$
$=E\biggl{[}\int_{t}^{t+\Delta t}[\mathcal{Q}_{\boldsymbol{u}}(t^{\prime})]_{ik}[\mathcal{Q}_{\boldsymbol{u}}(t^{\prime})]_{jk}\mathrm{d}t^{\prime}\biggl{|}\boldsymbol{u}(t)\biggr{]}$
$=E\biggl{[}\int_{t}^{t+\Delta t}[\mathcal{Q}_{\boldsymbol{u}}(t^{\prime})]_{ij}\mathrm{d}t^{\prime}\biggl{|}\boldsymbol{u}(t)\biggr{]}$
$\simeq[\mathcal{Q}_{\boldsymbol{u}}(t)]_{ij}\Delta t\,;\quad\Delta t\to 0\,;$ (57)
where the summation convention has been applied. Here, we used (i) the properties of the Itō integral, in particular, that the quadratic variation of $\boldsymbol{W}(t)$ is $\mathrm{d}W_{k}(t)\cdot\mathrm{d}W_{l}(t)=\delta_{kl}\mathrm{d}t$ in mnemonic form, (ii) that $\mathcal{Q}_{\boldsymbol{u}}$ is a symmetric matrix with $\mathcal{Q}_{\boldsymbol{u}}^{2}=\mathcal{Q}_{\boldsymbol{u}}$, and (iii) the continuity of the integrand. Expressing $\boldsymbol{W}^{\perp}(t)$ in terms of $\delta\boldsymbol{\omega}_{\perp}(t)$, one sees that Eq. 57 corresponds to Eq. 50b.